OpenStudy (anonymous):

One man starts out from a city and walks north at the rate of 3 kilometers per hour. Three hours later, a second man starts from the same place and walks the same route at 4 kilometers per hour. How far will he have to go before he overtakes the first man?

OpenStudy (lgbasallote):

let t = man A's time t + 3 = man B's time when he overtakes the guy they would be at the same distance from the starting point d = speed x time d of man A = 3t d of man B = 4(t-1) like i said if they overtake then distances are the same 3t = 4(t-1) solve for t

OpenStudy (lgbasallote):

oops i did something wrong idk why

OpenStudy (lgbasallote):

d of man B = 4(t+3) like i said if they overtake then distances are the same 3t = 4(t+3) solve for t

OpenStudy (lgbasallote):

@ganeshie8 did i do it right? im smelling negative time -_-

OpenStudy (anonymous):

hmmm im not getting you

OpenStudy (anonymous):

@ganeshie8 do you understand this one? i need your help.

ganeshie8 (ganeshie8):

4t-9 = 3t solve for t

OpenStudy (lgbasallote):

9?

ganeshie8 (ganeshie8):

wait.. checking im also haunted by -ves... lol

OpenStudy (anonymous):

yes 9 there but then what? and to be honest i dont understand why you arranged that as you did...

OpenStudy (lgbasallote):

lol =))) traumatized :))

OpenStudy (anonymous):

haunted???traumatized??? what are yall talking about?

ganeshie8 (ganeshie8):

ok.. you unerstand straight lines in slope-intercept form ?

OpenStudy (anonymous):

dont think so

OpenStudy (anonymous):

you still working on it?

ganeshie8 (ganeshie8):

ok.. as @lgbasallote said, distance is same here right. distance travelled by First man in time t = 3t in 3 hours he would have travelled 9 km so 9km is the distance second man lags when he starts his run. he is running at 4t speed. so at any point he will be covering (4t-9km) distance so far does it make sense ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please continue

ganeshie8 (ganeshie8):

so, now we can equate the distance and get the value of t. distance travelled by first man : 3t distance travelled by second man : 4t-9 4t-9 = 3t => t = 9 hours so, after 9 hours, both guys would have travelled the same distance. but question is asking about the distance second man has to travel before he overtakes first man, right ?

OpenStudy (anonymous):

yah

OpenStudy (saranya):

You can do this in two ways, do it in one way and check it using the second way. The first is to use the speed/distance/time formula and the second is to use a velocity-time graph and use the areas.

OpenStudy (vishweshshrimali5):

I can surely help What we have to do is just to use the concept of relative velocity

OpenStudy (vishweshshrimali5):

FIrst of all, Find out what distance does first person travel in those 3 hours

OpenStudy (vishweshshrimali5):

Now second person has to travel this distance + some more distance. Now relative velocity of second person wrt first person = 4-3 km/h = 1 km/h Now he has to travel a distance equal to distance travelled by first person in 3 h and that also with a relative speed of 1 km/h. From here find out the time

OpenStudy (anonymous):

hmm confusing

OpenStudy (vishweshshrimali5):

u start solving step by step . I am here to help u

OpenStudy (anonymous):

ok well all i know is 3k per hour and 3 hours later equals a 9kmperhour advantage.

OpenStudy (vishweshshrimali5):

Only 9 km advantage in this 3 hour interval. Now use the concept of relative velocity First person let A is moving with 3 km/h Second person say B is moving with 4 km/h So B has an advantage of 1 km/h Right ? This is the relative velocity of B w.r.t A

OpenStudy (vishweshshrimali5):

Got this ?

OpenStudy (anonymous):

yes he can over take him because he goes a bit faster got that

OpenStudy (vishweshshrimali5):

Yes gud Now overall B has to travel 9 km at a speed of 1 km/h Right, I am taking relative speed here as that will help me in assuming that A is at rest. Now, time taken by B to travel these 9 km = 9 h Got this ?

OpenStudy (anonymous):

yes to eliminate that advantage

OpenStudy (vishweshshrimali5):

Now what distance B actually travelled during these 9 hours = 9 * 4 = 36 km

OpenStudy (vishweshshrimali5):

I hope I did that correctly ?

OpenStudy (anonymous):

then the other guy went 9x3 = 27 + the 9 advantage =36 so they both tie at 36?

OpenStudy (vishweshshrimali5):

Of course

OpenStudy (anonymous):

but a tie is not an overtake

OpenStudy (vishweshshrimali5):

It is not that dear, If you have tied it than it will not even take a second for B to overtake A as he has an advantage of 1 km/h speed so that wouldn't be right to tell when B will overtake

OpenStudy (anonymous):

wouldnt?

OpenStudy (vishweshshrimali5):

Yes ofcourse

OpenStudy (anonymous):

hmm i think i lost you

OpenStudy (vishweshshrimali5):

What happened

OpenStudy (anonymous):

i dont get the tie/overtake thing

OpenStudy (vishweshshrimali5):

See If I say u that me and u r having a race I was ahead of u when u started but u started with a higher speed A time came when I and U were at same point but as you are having more speed than me U will easily overtake me even without taking a second after the tie point Got this ? So, that's why in such case we calculate the tie point as obviouslly after a very small time u will be ahead of me

OpenStudy (anonymous):

hmm but to be technical the the answer would be a tiny bit over 36

OpenStudy (vishweshshrimali5):

Yes but that tiny but will be very very very close to 0

OpenStudy (anonymous):

yes i see so 36 then

OpenStudy (vishweshshrimali5):

yes

OpenStudy (vishweshshrimali5):

did that help u ?

OpenStudy (anonymous):

yes very much thank you i may message you one like this but a bit harder